50. A rock is thrown from the top of a tall building. The distance in feet, between the rock and the ground t seconds after it is thrown is given by d= -16t^2-2t+524. How long after the rock is thrown is it 321 feet from the ground?

Answer by checkley75(3666)

321=-16t^2-2t+524
-16t^2-2t+524-321=0
-16t^2-2t+203=0
using the quadratic equation we get:
t=(2+-sqrt[-2^2-4*-16*203])/2*-16
t=(2+-sqrt[4+12992])/-32
t=(2+-sqrt12996)/-32
t=(2+-114)/-32
t=(2+114)/-32
t=116/-32
t=-3.625 not the answer.
x=(2-114)/32
t=-112/-32
t=3.5 seconds is the answer.